Research Index / HERCULES / Schedule of previous weeks / 2004
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC | PETAWATT |
|
01/05 Monday |
Regen in full power , SPIDER alignment |
second compressor alignment |
|
yellow channel pump alignment |
|
01/06 Tuesday |
Regen in full power, second order autocorrelator alignment, pulse width measurements |
|
yellow channel pump alignment | |
|
01/07 Wednesday |
Regen in full power, cleaner on, 4-pass amplifier on, |
Firing with 10 TW into solid target, proton beam profile measurements |
yellow channel pump alignment |
|
|
01/08 Thursday |
Regen in full power, cleaner on, 4-pass amplifier on, |
Firing with 10 TW into solid target, proton beam profile measurements |
yellow channel pump alignment |
|
|
01/09 Friday |
Regen in full power, cleaner on, 4-pass amplifier on, |
Firing with 10 TW into solid target, proton beam divergence measurements |
yellow channel pump alignment |
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC | PETAWATT |
|
01/12 Monday |
Regen in full power, cleaner on, 4-pass amplifier on, |
Firing with 10 TW into solid target, proton beam profile measurements |
|
yellow channel pump alignment |
|
01/13 Tuesday |
Regen in full power, SPIDER alignment, vibration problems |
|
yellow channel pump alignment | |
|
01/14 Wednesday |
Regen in full power, looking for the source of instability, SPIDER alignment |
|
|
|
|
01/15 Thursday |
Regen in diagnostic regime. Strong vibrations. |
yellow channel pump alignment |
||
|
01/16 Friday |
Regen in diagnostic regime. the group was looking for the source of vibrations.
|
|
yellow channel pump alignment |
From 19 up to 25 of January our group was looking for the source of laser instability. The reason had been found out on Sunday (January 25), it was a shutter. The current which goes through the shutter's electromagnet heats up metal parts and forms streams of warm air. The fluctuations of refraction coefficient of warm air stream brings the instability.
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC | PETAWATT |
|
01/26 Monday |
Regen in diagnostic power |
alignment |
|
|
|
01/27 Tuesday |
Regen in diagnostic power |
alignment |
||
|
01/28 Wednesday |
Regen in full power, cleaner on, 4-pass amplifier on. |
Firing with 10 TW into solid target, proton beam profile measurements |
||
|
01/29 Thursday |
Regen in diagnostic power.
Pumping lamps has been changed in 100TW stage |
imiging diagnostic alignment | ||
|
01/30 Friday |
Regen in diagnostic power Pumping channel on 100TW stage alignment. |
imiging diagnostic alignment
|
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC | PETAWATT |
|
02/02 Monday |
Regen in diagnostic power |
parabola alignment |
|
|
|
02/03 Tuesday |
Regen in diagnostic power |
parabola alignment |
||
|
02/04 Wednesday |
Regen in full power, cleaner on, 4-pass amplifier on. |
Firing with 10 TW into solid target, proton beam profile measurements |
||
|
02/05 Thursday |
Regen in diagnostic power |
alignment |
||
|
02/06 Friday |
Regen in full power, cleaner on, 4-pass amplifier on. |
Firing with 10 TW into solid target, proton beam profile measurements |
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC | PETAWATT |
|
02/09 Monday |
Regen in full power, cleaner on, 4-pass amplifier on, 2-pass amplifier on |
Firing with 30 TW into solid target, proton energy measurements |
|
|
|
02/10 Tuesday |
Regen in diagnostic power, 4-pass amplifier alignment to improve beam profile |
target alignment |
||
|
02/11 Wednesday |
Regen in diagnostic power |
alignment |
||
|
02/12 Thursday |
Regen in full power, cleaner on, 4-pass amplifier on, 2-pass amplifier on |
Firing with 30 TW into solid target, proton energy measurements |
||
|
02/13 Friday |
Regen in full power, cleaner on, 4-pass amplifier on, 2-pass amplifier on |
alignment |
The week Monday 02/15 - Friday 02/20 has been devoted to oscillator, regen , 4-pass and SPIDER
alignments. In the field of above threshold ionization experiment experiment the experimental chamber stabilization and vacuum level improvement has been done.February 2004
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC | PETAWATT |
|
02/23 Monday |
Oscillator chiller problems |
vacuum improvement |
|
|
|
02/24 Tuesday |
Oscillator and regen alignment. Regen in diagnostic power. |
alignment |
||
|
02/25 Wednesday |
Regen in diagnostic power, regen in full power. |
alignment, vacuum problems |
||
|
02/26 Thursday |
Regen in full power |
single atom ionization by above threshold ionization experiment |
||
|
02/27 Friday |
Regen in full power, cleaner on, 4-pass amplifier on. |
solid target ionization. proton spectrum measurement |
In March the highest intensity of 0.8´1022 W/cm2 was achieved by using a large NA paraboloid corrected by a deformable mirror. Focused intensity with high spatial resolution is calculated from the precise wavefront characterization.
In April and May the experiments on firing with above threshold ionization experiment into rarified gases has been done. One more channel in petawatt pump has been created and aligned.
May 2004
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC | PETAWATT |
|
05/17 Monday |
Regen in diagnostic power |
setup modification and alignment |
|
Blue pump channel alignment |
|
05/18 Tuesday |
Regen in diagnostic power |
setup modification and alignment |
Blue pump channel alignment |
|
|
05/19 Wednesday |
Regen in diagnostic power |
setup modification and alignment |
Frequency doubling crystal installation in blue channel |
|
|
05/20 Thursday |
Regen in full power |
setup modification and alignment |
Second harmonics power measurement in blue channel | |
|
05/21 Friday |
Regen in full power |
setup modification and alignment | ||
|
05/22 Saturday |
Regen in full power |
setup modification and alignment |
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC | PETAWATT |
|
05/22 Sunday |
Regen in full power. |
setup modification and alignment | ||
|
05/23 Monday |
Regen in diagnostic power |
setup modification and alignment |
|
Blue pump channel alignment |
|
05/24 Tuesday |
Regen in diagnostic power |
setup modification and alignment |
Blue pump channel alignment |
|
|
05/25 Wednesday |
Regen in diagnostic power |
setup modification and alignment |
Frequency doubling crystal installation in blue channel |
|
|
05/26 Thursday |
Regen in full power |
setup modification and alignment |
Second harmonics power measurement in blue channel | |
|
05/27 Friday |
Regen in full power |
setup modification and alignment |
||
|
05/28 Saturday |
Regen in full power |
setup modification and alignment |
June 2004
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC | PETAWATT |
|
05/30 Sunday |
Regen in full power | setup modification and alignment | ||
|
05/31 Monday |
Regen in full power, 4-pass on |
Firing with 10 TW into rarified neon |
|
glass laser maintenance |
|
06/01 Tuesday |
Water shout down | glass laser maintenance | ||
|
06/02 Wednesday |
Regen in diagnostic power |
|
setup modification and alignment |
|
|
06/03 Thursday |
Regen in diagnostic power |
|
setup modification and alignment | |
|
06/04 Friday |
Water shout down |
|
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC | PETAWATT |
|
06/06 Sunday |
Regen in full power | setup modification and alignment | ||
|
06/07 Monday |
Regen in diagnostic power |
above threshold ionization experiment |
Alignment of system with F2 paraboloid – optimization of focal spot, positioning of supersonic nozzle and installation of diagnostics (top scatter, fluorescent screen and ICT) |
|
|
06/08 Tuesday |
Regen in full power, optimization of compressor grating to get shortest pulse, 4-pass on | setup modification and alignment |
Completion of experimental alignment, high power shots (1-10 TW) |
|
|
06/09 Wednesday |
Regen in full power |
|
Completion of experimental alignment. |
|
|
06/10 Thursday |
Regen in full power, chiller problems |
|
Completion of experimental alignment. Vacuum improvement. |
|
|
06/11 Friday |
Regen in full power, 4-pass on. |
|
The vacuum in LILAC chamber was not deep enough to shoot. |
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC | PETAWATT |
|
06/13 Monday |
Macho Light chilling system maintenance. Regen in diagnostic power |
|
Vacuum improvement |
|
|
06/14 Tuesday |
Regen in full power, 4-pass on |
Completion of experimental alignment, high power shots (1-10 TW) |
||
|
06/15 Wednesday |
Regen in full power, 4-pass on |
|
Completion of experimental alignment, high power shots (1-10 TW) |
|
|
06/16 Thursday |
Regen in full power, 4-pass on |
|
Completion of experimental alignment, high power shots (1-10 TW) |
|
|
06/17 Friday |
Optimization of compressor grating to get shortest pulse. |
|
Completion of experimental alignment |
|
DATE |
HERCULES | ABOVE THRESHOLD IONIZATION EXPERIMENT | LILAC (Sudeep Banerjee) | PETAWATT |
|
06/21 Monday |
Macho Light chilling system maintenance. Regen in diagnostic power |
|
Vacuum improvement |
|
|
06/22 Tuesday |
Regen in full power |
Completion of experimental alignment, |
||
|
06/23 Wednesday |
Regen in full
power, 4-pass on Chilling system problems |
|
Completion of experimental alignment, high power shots (1-10 TW) |
|
|
06/24 Thursday |
Macho Light chilling system maintenance. |
|
|
New YLF crystal has been installed in osscilator of Glass Laser |
|
06/25 Friday |
Regen in full power, 4-pass on |
|
High power shots (1-10 TW) |
Frequency doubling crystal installation for blue channel of Petawatt pump |
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC (Banerjee) |
PETAWATT |
|
06/28 Monday |
Regen in full power, 4-pass on |
|
Completion of experimental alignment, high power shots (1-10 TW) |
|
|
06/29 Tuesday |
Regen in full power, 4-pass on Two mirrors in 4-pass schematic has been burned, most probable reason is vibration. The mirrors are changed and entire system is realigned. |
|
|
Energy measurement schematic for blue channel has been created |
|
06/30 Wednesday |
Electronics problems with Dazzler, the Dazzler has been taken out of schematic and system realigned |
|
|
|
|
07/01 Thursday |
Regen in full power, 4-pass on, 2-pass on |
|
High power shots (10-40 TW) |
|
|
07/02 Friday |
Regen in full power, 4-pass on, 2-pass on |
|
High power shots (10-40) TW) |
|
After July 4 it is going to be a construction work around Gerstacker building. We do not plan any experiments until July 20.
Since July 4 up to September 3 a big construction of parking lot area had been taking place. Digging machines caused huge vibrations. High energy shots were impossible under such vibrations. Only one (August 10) time we could fire 30 TW for Sudeep Banerjee experiment. The rest of time has been devoted to Hercules laser maintenance, contrast improving, SPIDER system development and petawatt pump development. In frameworks of laser maintenance we changed a laser diode in oscillator pump laser, realigned regen exit, filled cryo pump with new gas, cleaned and placed back both Ti:Sap crystals from four- and two-pass amplifiers, realigned compressor, improved the cooling group of Macho Light laser (four-pass pump). The spectrometer from John Nees lab has been used to run SPIDER, the results of phase measurements are not very impressive, but the SPIDER was working and it was confirmed that laser pulse compression has been screwed up by vibrations. Second channel (blue) and entrance in third one (green) in petawatt pump has been built. For a contrast improvement we are going to send laser beam through the hollow fiber placed in the chamber with high Ar or Kr pressure and to use non-linear rotation of polarization. Now we already have good laser spot ( 100 fs pulses from Ti:Sap) distribution at the hollow fiber exit and work on high pressure chamber.
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC (Banerjee) |
PETAWATT |
|
09/06 Monday |
LABOR DAY |
|
|
|
|
09/07 Tuesday |
Regen in full power, Compressor chamber alignment , second order autocorrelation measurements has been done |
|
Imaging system alignment |
Energy measurement schematic for blue channel has been created |
|
09/08 Wednesday |
Regen in full power |
Fiber targets has been produced |
alignment |
|
|
09/09 Thursday |
Regen in full power, 4-pass on, 2-pass on |
|
High power shots (10-40 TW) |
|
|
09/10 Friday |
Regen in full power, 4-pass on, 2-pass on |
|
High power shots (10 TW), the pressure in LILAC chamber went up and 40 TW shots has been cancelled |
|
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC (Banerjee) |
PETAWATT |
|
09/13 Monday |
CryoTiger pump maintenance |
|
|
Head A in Glass laser has been taken out for reflector repair |
|
09/14 Tuesday |
Regen in full power, 4-pass on |
|
|
|
|
09/15 Wednesday |
Regen in full power |
Breakdown threshold experiment |
|
|
|
09/16 Thursday |
Regen in full power, SPIDER measurements |
|
|
|
|
09/17 Friday |
Regen in full power, SPIDER measurements, PRO pump problems |
|
|
|
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC (Banerjee) |
PETAWATT |
|
09/20 Monday |
Regen pump maintenance, Glass Laser reflector repair |
|
|
|
|
09/21 Tuesday |
Regen pump maintenance, Glass Laser reflector repair |
|
|
|
|
09/22 Wednesday |
Regen in diagnostic power |
alignment |
|
|
|
09/23 Thursday |
Regen in diagnostic power |
alignment |
|
|
|
09/24 Friday |
Regen in full power |
alignment |
|
|
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC (Banerjee) |
PETAWATT |
|
09/27 Monday |
Regen in full power |
Alignment |
|
Head A of Glass laser has been assembled and tested |
|
09/28 Tuesday |
Regen in full power, pointing stability system adjustment |
|
|
The head A has been installed in Glass laser system |
|
09/29 Wednesday |
Regen in full power, 4-pass on |
10-5 TW shots |
|
|
|
09/30 Thursday |
Regen in full power |
alignment |
|
Calibration of energy measurement schematic in Petawatt Blue channel |
|
10/01 Friday |
Regen in full power, 4-pass on |
10-5 TW shots |
|
Energy measurements in Blue channel |
|
DATE |
HERCULES |
ABOVE THRESHOLD IONIZATION EXPERIMENT |
LILAC (Banerjee) |
PETAWATT |
|
10/04 Monday |
Regen in full power, 4-pass on |
10 TW shots |
|
Energy measurements in Blue channel |
|
10/05 Tuesday |
Regen in full power, pointing stability system adjustment |
10 TW shots |
|
Energy measurements in Blue channel |
|
10/06 Wednesday |
Regen in full power, 4-pass on |
10 TW shots |
|
|
|
10/07 Thursday |
Regen in full power, 4-pass on |
10 TW shots |
|
Green channel alignment |
|
10/08 Friday |
Regen in full power, 4-pass on |
10 TW shots |
|
Green channel alignment |
| October 11, 2004 - December 23, 2004 |
The HERCULES laser was used by the group of Donald Umstadter and Sudeep Barenjee for gas jet target experiments. |
Research Index / Hercules / Schedule of previous weeks / 2004
|
Copyright © Center for Ultrafast Optical Science, University of Michigan |
|
Maintained
by: |